16. A drop of water has a volume of 50 uL. How many water molecules does it contain? The density of water is 1.0 g/mL. A. 1.6Please show all steps and explain, as I need to understand for an exam. Thx!
Answer #1

Volume = 50 uL

= 0.050 mL

USE:

Mass of H2O = density * volume

= 1.0 g/mL * 0.050 mL

= 0.050 g

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

mass(H2O)= 0.050 g

use:

number of mol of H2O,

n = mass of H2O/molar mass of H2O

=(5*10^-2 g)/(18.02 g/mol)

= 2.775*10^-3 mol

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 2.775*10^-3 * 6.022*10^23 molecules

number of molecules = 1.67*10^21 molecules

Answer: A